# Hierarchical Universal Value Function Approximators

### v1

@@ 1 Introduction
Value functions ($V(s)$ and $Q(s,a)$) are key to reinforcement learning algorithms (Sutton & Barto, 1998). They represent how important it is for the agent to be in a certain state and take certain actions in a given state. Using an optimal value function for a particular goal allows us to recreate the optimal policy the agent must follow to maximize the discounted return for that goal. Hierarchical reinforcement learning extends this to temporal abstraction (Sutton et al., 1999; Precup, 2000) by using a hierarchy of two or more levels of value functions to create a collection of useful skills or behaviors (Silva et al., 2012) that solve independent tasks or subgoals. A meta-value function selects the appropriate skill in the given state while a lower-level value function drives the policy that executes the actions the agent takes when exploiting the chosen skill. This is beneficial for scaling and planning.
@@
We extend UVFAs to hierarchical reinforcement learning, or temporal abstraction, using the options framework. Since there are two or more levels of value functions in a hierarchy, we must learn representations (both parameterized and non-parameterized) of all the levels of the hierarchy to results in a final state of value functions equal in number to the hierarchy levels. The challenge lies in extending the concept of UVFAs from two or three dimensions (for the state and action value functions respectively) to n-dimensions. Our method is novel because it shows that there is underlying structure in the states, goals, options and actions that results in a universal representation of the hierarchy. This universal representation results in universal temporally abstract behaviors or skills that has zero-shot generalization to unseen goals. We show that this representation outperforms corresponding UVFAs.
@@ 2 Background
Solving for the optimal action-value function results allows one to deduce a greedy optimal policy. For discrete MDPs there is at least one optimal policy that is greedy with respect to its action-value function.
@@
The Options Framework (Sutton et al., 1999; Precup, 2000) formalizes the concept of temporal abstractions in reinforcement learning, or hierarchical reinforcement learning, by using semi-MDPs. A Markovian option $\omega\in\Omega$ is a tuple $(\mathcal{I}_{\omega},\pi_{\omega},\beta_{\omega})$ where $\mathcal{I}_{\omega}\in\mathcal{S}$ is its initiation set, $\pi_{\omega}$ is its intra-option policy, and $\beta_{\omega}:\mathcal{S}\to[0,1]$ is its termination function. We assume that $\forall s\in\mathcal{S}$, $\forall\omega\in\Omega$: $s\in\mathcal{I}_{\omega}$–all options are available everywhere. An option $\omega$ is picked from the initiation set $\mathcal{I}_{\omega}$ using the policy-over-options $\pi_{\Omega}(s,o)$ and follows its intra-option policy $\pi_{\omega}(s,a)$ until it terminates by $\beta_{\omega}(s)$ which determines the probability of a given option terminating in a given state.
This is a call-and-return model where the option initiation is determined by the policy-over-options (meta-policy) $\pi_{\Omega}(\omega|s)$ that gives us the probability of selecting option $\omega$ in state $s$ and is driven by the option-value function $Q_{\Omega}(s,\omega):\mathcal{S}\times\Omega\to\mathbb{R}$. Once the option is selected, it performs actions using the intra-option policy $\pi_{U}(a|s,\omega)$ which is driven by its intra-option value function $Q_{U}(s,o,a):\mathcal{S}\times\Omega\times\mathcal{A\to\mathbb{R}}$. This continues until the option terminates, upon which the process restarts in the then-current state. The two sets of value functions are related as
$Q_{\Omega}(s,\omega)=\sum_{a}\pi_{\omega,\theta}\left(a\;\middle|\;s\right)Q_{U}(s,\omega,a)\enspace,$ (1)
$Q_{U}(s,\omega,a)=r(s,a)+\gamma\sum_{s^{\prime}}P\left(s^{\prime}\;\middle|\;s,a\right)U(\omega,s^{\prime})\enspace.$ (2)
@@
$U(\omega,s^{\prime})=(1-\beta_{\omega,\vartheta}(s^{\prime}))Q_{\Omega}(s^{\prime},\omega)+\beta_{\omega,\vartheta}(s^{\prime})V_{\Omega}(s^{\prime})\enspace$
is the value of executing $\omega$ upon entering a state $s^{\prime}$.
@@ 3 Hierarchical Universal Value Function Approximators
  • The intra-option H-UVFA $Q_{U}(s,o,a,g;\eta)$ can be split into three sets of embeddings: $\phi:\mathcal{S}\to\mathbb{R}^{n}$, $\psi:\mathcal{G}\to\mathbb{R}^{n}$, $\chi:\Omega\to\mathbb{R}^{n}$, and $\delta:\mathcal{A}\to\mathbb{R}^{n}$
@@ 4.2 Supervised Learning Experiments
Results: We construct H-UVFAs using a combination of 25 goals. Figure 2 plots the values and greedy actions of the H-UVFAs on a subset of the goals used during training along with the ground truth values corresponding to the option-value and intra-option value functions. In each state, the greedy option is picked and the corresponding value and greedy action are plotted. We find that an embedding rank of 50 is sufficient to recreate the ground truth matrix accurately. We plot the entire decomposed matrices and the corresponding recreated H-UVFAs for $Q_{\Omega}$ and $Q_{U}$ in Appendix B.
Interestingly, we see that the recreation of the value function with a neural network allows for more continuous changes in the values—as we move further away from the goal the value of the state decreases gradually as compared to the more sudden changes in value functions seen in the discrete and tabular ground truth. This behavior better represents a value function and policy and allows for better behavior in longer trajectories where more optimal actions are picked further away from the goal as seen in Figure 2.
@@ 5.2.2 Learning with Bootstrapping
Similar to UVFAs, it is possible to learn H-UVFAs directly using bootstrapping following value function updates::
@@ 6 Comparison with UVFAs
7 Ongoing Work
We are currently working on extending the results to a more complex domain in the Arcade Learning Environment (Bellemare et al., 2013), and will update the paper with further results.
@@ Appendix A Experiment Details
Appendix B Factorized Matrices and Recreations
@@
Figure 9: The recreation of the meta-policy and option-policies for a single goal with the embeddings trained on 25 goals. The ground truth is in blue with the recreations in red.

### v2

@@ 1 Introduction
Value functions ($V(s)$ and $Q(s,a)$) are key to reinforcement learning algorithms (Sutton & Barto, 1998). They represent how important it is for the agent to be in a certain state and take certain actions in a given state. Using an optimal value function for a particular goal allows us to re-create the optimal policy the agent must follow to maximize the discounted return for that goal. Hierarchical reinforcement learning extends this to temporal abstraction (Sutton et al., 1999; Precup, 2000) by using a hierarchy of two or more levels of value functions to create a collection of useful skills or behaviors (Silva et al., 2012) that solve independent tasks or subgoals. A meta-value function selects the appropriate skill in the given state while a lower-level value function drives the policy that executes the actions the agent takes when exploiting the chosen skill. This is beneficial for scaling and planning.
@@
We extend UVFAs to hierarchical reinforcement learning, or temporal abstraction, using the options framework. Since there are two or more levels of value functions in a hierarchy, we must learn representations (both parameterized and non-parameterized) of all the levels of the hierarchy to result in a final state of value functions equal in number to the hierarchy levels. The challenge lies in extending the concept of UVFAs from two or three dimensions (for the state and action value functions respectively) to n-dimensions. Our method is novel because it shows that there is underlying structure in the states, goals, options and actions that results in a universal representation of the hierarchy. This universal representation results in universal temporally abstract behaviors or skills that has zero-shot generalization to unseen goals. We show that this representation outperforms corresponding UVFAs.
@@ 2 Background
Solving for the optimal action-value function allows one to deduce a greedy optimal policy. For discrete MDPs there is at least one optimal policy that is greedy with respect to its action-value function.
@@
The Options Framework (Sutton et al., 1999; Precup, 2000) formalizes the concept of temporal abstractions in reinforcement learning, or hierarchical reinforcement learning, by using semi-MDPs. A Markovian option $o\in\Omega$ is a tuple $(\mathcal{I}_{o},\pi_{o},\beta_{o})$ where $\mathcal{I}_{o}\in\mathcal{S}$ is its initiation set, $\pi_{o}$ is its intra-option policy, and $\beta_{o}:\mathcal{S}\to[0,1]$ is its termination function. We assume that $\forall s\in\mathcal{S}$, $\forall o\in\Omega$: $s\in\mathcal{I}_{o}$–all options are available everywhere. An option $o$ is picked from the initiation set $\mathcal{I}_{o}$ using the policy-over-options $\pi_{\Omega}(s,o)$ and follows its intra-option policy $\pi_{o}(s,a)$ until it terminates by $\beta_{o}(s)$ which determines the probability of a given option terminating in a given state.
This is a call-and-return model where the option initiation is determined by the policy-over-options (meta-policy) $\pi_{\Omega}(o|s)$ that gives us the probability of selecting option $o$ in state $s$ and is driven by the option-value function $Q_{\Omega}(s,o):\mathcal{S}\times\Omega\to\mathbb{R}$. Once the option is selected, it performs actions using the intra-option policy $\pi_{U}(a|s,o)$ which is driven by its intra-option value function $Q_{U}(s,o,a):\mathcal{S}\times\Omega\times\mathcal{A\to\mathbb{R}}$. This continues until the option terminates, upon which the process restarts in the then-current state. The two sets of value functions are related as
$Q_{\Omega}(s,o)=\sum_{a}\pi_{o,\theta}\left(a\;\middle|\;s\right)Q_{U}(s,o,a)\enspace,$ (1)
$Q_{U}(s,o,a)=r(s,a)+\gamma\sum_{s^{\prime}}P\left(s^{\prime}\;\middle|\;s,a\right)U(o,s^{\prime})\enspace.$ (2)
@@
$U(o,s^{\prime})=(1-\beta_{o,\vartheta}(s^{\prime}))Q_{\Omega}(s^{\prime},o)+\beta_{o,\vartheta}(s^{\prime})V_{\Omega}(s^{\prime})\enspace$
is the value of executing $o$ upon entering a state $s^{\prime}$.
@@ 3 Hierarchical Universal Value Function Approximators
  • The intra-option H-UVFA $Q_{U}(s,o,a,g;\eta)$ can be split into four sets of embeddings: $\phi:\mathcal{S}\to\mathbb{R}^{n}$, $\psi:\mathcal{G}\to\mathbb{R}^{n}$, $\chi:\Omega\to\mathbb{R}^{n}$, and $\delta:\mathcal{A}\to\mathbb{R}^{n}$
@@ 4.2 Supervised Learning Experiments
Results: We construct H-UVFAs using a combination of 25 goals. Figure 2 plots the values and greedy actions of the H-UVFAs on a subset of the goals used during training along with the ground truth values corresponding to the option-value and intra-option value functions. In each state, the greedy option is picked and the corresponding value and greedy action are plotted. We find that an embedding rank of 50 is sufficient to re-create the ground truth matrix accurately. We plot the entire decomposed matrices and the corresponding re-created H-UVFAs for $Q_{\Omega}$ and $Q_{U}$ in Appendix B.
Interestingly, we see that the re-creation of the value function with a neural network allows for more continuous changes in the values—as we move further away from the goal the value of the state decreases gradually as compared to the more sudden changes in value functions seen in the discrete and tabular ground truth. This behavior better represents a value function and policy and allows for better behavior in longer trajectories where more optimal actions are picked further away from the goal as seen in Figure 2.
@@ 5.2.1 Horde Generalization
Atari Experiments: Like Schaul et al. (2015), we scale our method for learning H-UVFAs to the Arcade Learning Environment (Bellemare et al., 2013) with the Ms. Pacman domain. The training goal set for the agents in the Horde consists of 30 pellets selected randomly throughout the space; the remaining 120 pellets are test goals used to measure generalization of the H-UVFAs. Each pellet is the only visible goal to the agent with all the other pellets being masked. The reward function is modified to provide a positive reward only upon reaching the correct pellet, with no reward being provided for any other pellet. Further experiment details can be found in Appendix A.
Figure 7 plots the distribution of the training and test pellets as well as the ground truth and H-UVFA retrieved value functions for three test goals. The H-UVFAs learned using the smaller Horde of agents can accurately approximate the value-functions of the unseen pellets. The policy learned from following the H-UVFAs is comparable to the ground truth policy due to the accuracy in the value function re-creation. H-UVFAs are able to exploit the underlying structure in the states, goals, options, and actions in higher dimensional spaces.
Figure 7: H-UVFA results on higher dimensional Ms. Pacman Domain. The top row plots the training and test pellet distribution in green and yellow respectively. The center and bottom rows depict the ground truth and H-UVFA re-created value functions respectively on test goals. Higher values are indicated in red, lower-values are indicated in blue, and the test goal is indicated in green. H-UVFAs can exploit the underlying structure in states, goals, options, and actions in higher dimensions as seen by their generalization when re-creating value functions for unseen goals.
@@ 5.2.2 Learning with Bootstrapping
Similar to UVFAs, it is possible to learn H-UVFAs directly using bootstrapping following value function updates:
@@ Appendix A Experiment Details
Atari: The Ms. Pacman experiments are performed on a modified environment where each pellet represents a goal. The training set consists of 30 pellets and the test set for generalization consists of the remaining 120 pellets. Ms. Pacman is initialized at a random position at the start of each episode. A PyTorch implementation of the Option-Critic architecture is used to train each agent in the Horde. The inputs are grayscale images processed with a convolutional network.
Appendix B Factorized Matrices and Re-creations
@@
Figure 10: The re-creation of the meta-policy and option-policies for a single goal with the embeddings trained on 25 goals. The ground truth is in blue with the re-creations in red.